OpenStudy (anonymous):

3x-2y=-12 x=2/3y-4 Solve each system by the substation method. NO SOLUTION OR AN INFINITE NUMBER

hero (hero):

\(3x - 2y = -12\) \(x = \dfrac{2}{3y - 4}\)

hero (hero):

Solve: \(3 \left(\dfrac{2}{3y - 4} \right) - 2y = -12\)